OpenStudy (anonymous):

-EXTREMELY URGENT- *Simplifications are accepted if no other way is found* I need to DETERMINE the FORCE exerted by an ELECTROMAGNET to an IRON SPHERE (or any easy shape)- - I'll place an iron sphere on a *scale* and put an electromagnet *directly* above it (some distance) - I don't yet know how high above it. How can I compare the mass that is lifted directly up with respect to field strength generated by the electromagnet in theory without having to practically test? Thanks in advance, this is extremely important. Any ideas are welcome.

OpenStudy (anonymous):

. I need the simplest experiment where I can compare the exertion of magnetic force to a physical phenomena resulting from this - other experiment setups are therefore welcome as well. Here you have something on the sphere possibility but I don't understand some parts of it. My physics understanding is that of a 12th grader. https://sbms.bnl.gov/sbmsearch/subjarea/118/1u04e011.pdf

OpenStudy (ghazi):

see...basically the ability of lifting certain mass by some magnetic force = m*g...also it depends on the capacity of magnet......and you can obtain relation between varying magnetic field and lifting of mass by a set of data ....and using curve fitting method...i am sure you'll get a direct relation between them

OpenStudy (anonymous):

I need a theory that an experiment that I will do supports. I know a higher current applied to an electromagnet increases the weight of iron that it attracts, but I need a solid theory to support this.

OpenStudy (ghazi):

well....can you form a set of data between varying magnetic field and weight lifted....if you can make it...then after that ..use curve fitting relation ...you'll have a direct function that relates magnetic field and amount of weight lifted....you can do it easily...

OpenStudy (anonymous):

I know varying magnetic field lifts masses to different extents, but what is a formal theory of this with equations that I can use to *predict* that by forming a set of data between varying magnetic field strength and weight lifted I will get a curve expressing how change in magnetic field strength changes the mass lifted. Equations, definitions, rules, how can I research this?

OpenStudy (theeric):

Additionally, I have some equations that could possibly combined into one experiment. 1. Magnetic field magnitude of a flat coil with N turns, area A, and radius R, at a distance of z away from the face of the loop. This is for a point, however. \[B=(\frac{\mu_0}{2\pi})*\frac{NiA}{(R^2+z^2)^{\frac{3}{2}}}\] 2. Force of a magnetic field on a current-carrying wire (and finally a non-trivial link) http://hyperphysics.phy-astr.gsu.edu/hbase/magnetic/forwir2.html 3. Magnetic force between two straight, parallel, current-carrying wires (wire a and wire b). \[F_{b.due.to.a} = (\frac{\mu_0}{2\pi})*(\frac{i_a*i_b}{d})*L_b\] \[\mu_0 = 4\pi*10^{-7}\] in Volt*second/Ampere/meter or several other units ( http://en.wikipedia.org/wiki/Magnetic_constant - and you can mouse over blue colored units to see what they are). \[L_b\] is the length of wire b. \[i_a\] is the current through wire a. \[i_b\] is the current through wire b. \[d\] is the distance between the wires. Also, \[F_{a.due.to.b} = (\frac{\mu_0}{2\pi})*(\frac{i_a*i_b}{d})*L_a\]

OpenStudy (theeric):

I do realize this doesn't pertain well to the iron, sphere, sorry.

OpenStudy (anonymous):

You really know a lot! "You still need an experiment in which you compare measured results to expected (theoretical) results, right?" Yes. If I want to get an electromagnet that will work with your theory what would it be? Or is what you said just a statement of a regular electromagnet's properties? I don't know how to search based on the information given so it would be cool if you could clarify the electromagnet properties. Secondly No electromagnet is uniform in practice - is this negligible or do you have an idea of how to calculate the difference that real life experiment with non-uniform magnetic field makes? Then, "This is for a point, however" which is I guess why your theory/equations cannot be directly applied for iron sphere? Note now! The object in question, atm iron sphere, can be changed to anything you seem fit, all I need is an experiment : ) You can vary the shape and size - as long as the magnetic field strength applied to it and the mass of the object being attracted by the field suit a doable real-life experiment. What would you have in mind with regard to this? Just.. Immense thanks to you for taking the time for what you've done! Tomorrow (In about 20 hours) I need to give my supervisor a short presentation of what I'm doing so I would have to have an experiment setup figured out along with the background investigation that is used in my essay to justify that the experiment will work. If you have the time, get back to me asap.

OpenStudy (theeric):

The iron sphere is complicated because iron is a magnetic material, and I don't understand magnetization. I've never worked with magnetic materials, and they generate they're own magnetic field. From what I gather, this results from the external magnetic field source causing changes in the material. By changes, I mean that formerly randomized portions with internal magnetic fields are increased or decreased to match the external field. But, I don't think that's what needs to be calculated, because some other equations are used to describe the general trend of the magnetic fields produced. Also, that formula for a point only works for points along the axis which the solenoid is around. What I did in class once was derive this formula with calculus - calculating the field at that point from every point along a coil. And then, if it's like electric fields (I think it is) then magnetic field vectors can add or subtract. So multiple loops just adds or subtracts to the multitude of the magnetic field at all points. That's why, with a flat coil (theoretical, but not too far off in certain circumstances), you multiply the magnetic field strength with the number of loops.

OpenStudy (theeric):

And I would think the magnetic field within a solenoid would be nearly uniform with exception to area around the edges. Hopefully this can be considered negligable. In my response with the 3 theories, #2 and #3 were for straight wires, and the best match for the first theory is many insulated coils wrapped together like a wreath. Anyone feel free to correct me if necessary!

OpenStudy (anonymous):

With what material, shape and size (with enough mass to be measured though) would your theory work so that I could use it?

OpenStudy (theeric):

I don't know if I can get those equations to work for an iron sphere! I have to look at magnetism again. But you could possibly get the magnetization or something else by measuring how the iron sphere reacts in the magnetic field of a solenoid.